OpenStudy (anonymous):

PLEASE HELPPP!! 3x+y=4 x+y=2

OpenStudy (amorfide):

you can use many methods easiest for this would be substitution or elimination elimination would be to have the same coefficient of an unknown x or y since you have the same coefficient (1) for y we will look at the sign, whether positive or negative the idea is to eliminate the unknown you have positive 1 y in both equations, to eliminate and to make it 0 y you need to subtract so you can subtract the equations from one another so you get 3x+y-(x+y)=4-2

OpenStudy (anonymous):

can you help me work out one of the equations

OpenStudy (amorfide):

okay so since we are using elimination method we will subtract the equations to get 3x+y-(x+y)=4-2 which gives 2x=2 work out x substitute x into one of the original equations to get y

OpenStudy (anonymous):

x=1 and y=1